OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the approximate value of angle A. 22* 38* 142* 158*

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):



OpenStudy (anonymous):

again use the law of cosines

OpenStudy (anonymous):

\(a^2=c^2+b^2-2cb ~\cos A\)

OpenStudy (anonymous):

22*

OpenStudy (anonymous):

thnx:)

OpenStudy (anonymous):

np :)